# Image-to-Video Transfer Learning based on Image-Language Foundation Models: A Comprehensive Survey

### v1

@@ II Overview of Vision-Language Foundation Model
LLaVA-1.5 [128] extends LLaVA by replacing its image encoder with CLIP-ViT-L-336px and vision-language projection layer with MLPs, which achieve better multi-modality understanding. Moreover, the model also introduces a multi-patch partitioning strategy, which improves the efficiency of processing high-resolution images and facilitates dynamic-resolution inputs. It also extends the training data with academic benchmarks of varied vision tasks including visual question answer (e.g., OKVQA [129] and A-OKVQA [130]) and OCR-intensive tasks (e.g., OCRVQA[131] and TextCaps [132]). LLaVA-NeXT [133] improves LLaVA-1.5 by quadrupling input image resolution (up to 672×672) and supporting flexible aspect ratios inputs, enabling capturing finer visual detail. It enhances reasoning, OCR, and world knowledge through optimized visual instruction tuning while maintaining data efficiency with under 1M training samples. The model achieves superior performance across benchmarks, even surpassing Gemini Pro [^7] https://deepmind.google/models/gemini/pro/ in some tasks, while retaining LLaVA’s design and fast training scalability (e.g., 34B parameters trainable in nearly 1 day on 32 A100 GPUs).
@@
Discussion: All these foundation models not only provide a solid foundation for improving image understanding in the presence of natural language but also set the stage for their application in more complex video analysis tasks, where the integration of dynamic temporal information becomes essential. Among these image-based foundation models, MDETR and GroundingDINO demonstrate stronger capabilities in capturing fine-grained visual–textual correspondences, as they are explicitly trained with grounding supervision to align object-level regions with textual phrases. By contrast, models such as CLIP and BLIP are more adept at handling abstract or high-level visual–textual correspondences, largely due to large-scale contrastive or generative pretraining. Instruction-tuned models like LLaVA bridge the two extremes by leveraging large language models to reason over both detailed and abstract alignments, though often at the expense of precision in object-level grounding.
III Overview of Image-to-Video Transfer Learning
In this section, we summarize common transfer paradigms for adapting image-level knowledge from Image-Language Foundation Models (ILFM) to address video-text learning tasks. Depending on whether the vision-language features extracted by ILFM remain unchanged or are adapted, we categorize existing transfer learning methods into frozen features and modified features, which will be elaborated in the following subsections.
@@ III-A Frozen Features
Figure 4: Methods and architectures of transferring pre-trained image-text model to video domain via frozen features.
By treating video as sequence of image frames in temporal order, many existing methods directly leverage pre-trained ILFM to extract frame-wise cross-modal features and then use the frozen frame-level features to facilitate their image-to-video transfer learning. Representative lines of work include knowledge distillation, post-network tuning, and side-tuning.
@@ III-A1 Knowledge distillation
As shown in Figure 4(A), within the knowledge distillation strategy, ILFM is primarily employed as a teacher model to enhance the cross-modal alignment capabilities of image/text encoders in existing video-text understanding models. Various distillation objectives can be defined to facilitate targeted knowledge transfer across encoders, working in conjunction with the primary task loss. Although this paradigm has been widely adopted in early Referring Video Object Segmentation (RVOS) and Open-Vocabulary Multi-Object Tracking (OV-MOT) methods to introduce open-vocabulary capabilities from models like CLIP, it often overlooks the discrepancy between image-level and video-level representations. This oversight may introduce erroneous knowledge from ILFMs, consequently limiting the effectiveness of cross-modal video-text understanding.
@@ III-A2 Post Network Tuning
Many methods achieve temporal modeling by additionally introducing a post network to process the frozen features obtained for each selected frame (shown in Figure 4(B)). It is straightforward to construct the post network with temporal attention mechanism, enabling the information exchange among varied frames. However, this paradigm struggles to effectively capture dynamic temporal patterns from static features, resulting in limited performance. To more effectively harness the visual potential of frozen large-scale image models, researchers employ a Q-Former [19] architecture, utilizing a set of learnable queries to interact with features from all video frames. Through this approach, queries can dynamically capture temporal dependencies and motion patterns in videos, rather than augmenting image features with temporal relations. This not only reduces computational costs but also unlocks the potential of ILFM. Beyond general temporal modeling, many methods have incorporated more task-specific temporal modeling strategies, such as the Hungarian algorithm or Brownian bridge-based association. In summary, this paradigm relies on high-quality image-text features and carefully designed, task-specific temporal modeling architectures to achieve strong video understanding performance.
@@ III-A3 Side-Tuning
as illustrated in Figure 4(C), Side-tuning [8] introduces a learnable side network alongside the frozen image and text encoders of ILFM. It receives the layer-wise latent features extracted by ILFM encoders as input, facilitating hierarchical spatial modeling and temporal refinement in a coarse-to-fine manner. In practice, the side network adopts a lightweight design, incorporating both text-modulated spatial pooling and temporal modeling operators [59]. During training, gradients are confined within the side-blocks, ensuring both parameter and memory efficiency. Recent study also show that the system can be further enhanced by connecting side-blocks with the outputs of CLIP, which achieves better results on several downstream video-text understanding tasks [59].
Discussion: Frozen-feature-based image-to-video transfer learning frameworks offer notable advantages in adapting image-based foundation models to video–language understanding tasks. They maintain high parameter efficiency by freezing pretrained architecture and weights, thereby mitigating catastrophic forgetting of previously acquired knowledge. However, their performance may be limited by the inherent representational capacity of static image-level features and their imperfect alignment with video-specific task requirements, particularly when generalizing to novel scenarios.
III-B Modified Features
Instead of directly using foundation models to extract frozen features, many approaches perform image-to-video transfer learning by explicitly adjusting the backbone model architecture and parameters, which we classify as modified features. In the following, we will elaborate on the techniques involved in modified features including fully-finetune, finetune with extra models, finetune with partial tuning, finetune with adapter, finetune with LoRA, and finetune with prompt tuning, whose architectures are presented in Figure 5, respectively.
Figure 5: Methods and architectures of transferring pre-trained image-text model to video domain via modified features.
@@ III-B1 Full Fine-Tuning
It is straightforward to empower a pre-trained image-text foundation model with the ability of video-text understanding by building a temporal module on the top of each block of image-language foundation model and then simply finetuning all the parameters involved in the temporal modules together with the ILFM [68, 70, 71, 69]. Although simple and effective, these approaches indeed are less applicable in practice as they need a large-scale well-defined and annotated data and a large amount of computing power in the model training stage.
@@ III-B2 Partial Tuning
Partial tuning strikes a balance between full fine-tuning and more parameter-efficient methods like adapters or prompt tuning. This strategy involves updating only a specific subset of the pre-trained ILFM’s parameters, while freezing the rest of the model parameters. Recently, several methods [134, 135, 136] have introduced temporal attention modules into the original spatial attention modules, enabling ILFMs to acquire temporal–spatial modeling capability. Such tuning strategies strike a favorable balance between performance and computational cost, while still relying heavily on the inherent spatial capacity of the ILFM, as they do not involve full model training.
@@ III-B4 Fine-Tuning with Adapter
Adapter-based methods offer a parameter-efficient solution to adapting ILFMs for addressing video-language understanding tasks, which intends to insert a set of lightweight, trainable module (referred as adapters) into some intermediate layers of pre-trained ILFMs  [139, 140, 60]. The adapter module are typically defined as a stack of temporal attention and cross-frame interaction operators in practice. By keeping the parameters corresponding to ILFM frozen, adapter-based approaches can retain the ability of visual-semantic understanding gained in pre-trained ILFM while efficiently learn temporal patterns specific to video data through the newly introduced adapter. Recently, the adapter-based finetuning techniques are particularly attractive in scenarios with limited video annotations or computational budgets, and they are emerging as a prominent direction for addressing downstream tasks like video question answering, temporal video grounding, and video captioning.
@@ III-B5 Fine-Tuning with LoRA
As a representative parameter-efficient fine-tuning (PEFT) method, Low-Rank Adaptation (LoRA) [141] provides an alternative paradigm for ILFM adapting. Instead of directly learning the updating of model weights $\Delta W$, LoRA select to approximate the updated weight with two small, trainable low-rank matrices, i.e., $\Delta W=BA$, where $B\in\mathbb{R}^{d\times r}$, $A\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$. During fine-tuning, the original parameters of the ILFM remain frozen and locked, preserving the extensive knowledge acquired during pre-training. Only the injected low-rank matrices $A$ and $B$ are updated, which dramatically reduces the number of trainable parameters and the accompanying GPU memory footprint. This approach implicitly constrains the model’s exploration within a low-dimensional subspace, effectively mitigating the risk of overfitting, which is particularly beneficial when labeled video data is scarce. Consequently, LoRA has gained significant traction for efficiently adapting large-scale foundation models to various downstream video understanding scenarios, establishing itself as a cornerstone technique in resource-constrained environments. However, as LoRA itself is incapable of modeling temporal information, this paradigm relies on external temporal modeling modules or networks, such as those used in fully finetuning or post-network tuning.
@@ III-B6 Prompt Tuning
Prompt tuning has recently emerged as a highly parameter-efficient alternative for adapting ILFMs to video-related tasks [122, 123]. By freezing the backbone and introducing a small set of learnable prompts, the model is guided to capture temporal cues without modifying its core parameters. A common design is to share a single set of prompts across all frames to enforce temporal consistency, or to propagate prompts sequentially across frames so that information accumulates and evolves with motion. These temporal variants enable the frozen ILFM to move beyond static frame modeling and capture motion cues and cross-frame dependencies. This strategy achieves efficient spatio-temporal adaptation with low computational cost, though its reliance on frozen backbones may restrict fine-grained temporal reasoning compared to full or partial fine-tuning.
Discussion: The aforementioned image-to-video transfer learning methods primarily adapt existing image–text foundation models to address video–text understanding tasks through two main directions: enhancing the encoders of established video understanding frameworks or incorporating additional learnable temporal modules into image–language foundation models. Notably, these approaches are not mutually exclusive. Existing studies have shown that their integration can yield complementary benefits and further improve transfer performance. For example, fine-tuning with auxiliary models can be seamlessly combined with parameter-efficient strategies such as LoRA, facilitating the development of more advanced and robust video understanding models.
@@ IV Fine-grained Video-Text Understanding
In the following, we review methods that adapt existing pre-trained image-text foundation models to address specific video-text understanding tasks. Based on the required granularity of understanding, we categorize the video-text understanding tasks into fine-grained and coarse-grained classes, providing a clearer structural understanding and enabling researchers to identify methods most relevant to their research domains. We first introduce fine-grained video-text understanding, which requires precise localization of spatial regions (e.g., bounding boxes or segmentations), temporal intervals, or targeted editing of relevant video segments based on textual descriptions.
IV-A Multi-Object Tracking
Open Vocabulary Multi-Object Tracking (OV-MOT) and Referring Multi-Object Tracking (RMOT) are two multi-object tracking tasks extended from traditional MOT, in which the objects of interest are explicitly referred with natural language descriptions. Specifically, OV-MOT aims to track all the objects  [^8] In this setting, the object number could be more than one and vary for different samples.described by input text in an open-vocabulary manner, rather than relying on a predefined, closed set of categories. The performance of OV-MOT systems is measured through metrics including localization accuracy (LocA), association accuracy (AssocA), and classification accuracy (ClsA). In contrast, Referring Multi-Object Tracking (RMOT) takes a natural language description specifying the objects of interest as input, and aims to detect and track these objects within a given video. This task is challenging as it requires the model to jointly capture fine-grained visual appearance, motion patterns, and semantic language cues for maintaining persistent trajectories of referred entities.
@@ IV-A Multi-Object Tracking
Most of existing OV-MOT methods [107, 41, 32, 29, 28, 40, 31, 30]mainly follow a frozen feature paradigm to transfer useful knowledge from pre-trained CLIP and Grounding-DINO models. For instance, OVTrack [28] and OVTrack+ [29] develop knowledge distillation blocks to enhance the RoI features extracted by ResNet50 and FPN with pre-trained CLIP, Based on OVTrack, VOVTrack [30] further introduce object appearance self-consistency and spatial-appearance mutual-consistency losses to capture spatial relationships across frames, where object appearance/attribute information are extracted with CLIP. SLAck [31] also distills textual knowledge from CLIP into its semantic head and introduces a spatial-temporal object attention mechanism to jointly model semantic, location, and appearance cues. GLATrack [32] extracts semantic information about the target objects from the reference frame using pre-trained CLIP model, which is then employed to reduce the influence of erroneous features in key frames. OVTR [41] employs a post-network fine-tuning strategy, wherein a Transformer-based dual-branch decoder is introduced to process frozen features extracted by CLIP. These methods rely on coarse-grained image-text correspondences captured by CLIP, which restricts the model’s ability to capture the open-world objects with high dynamicity. MASA [107] introduces a plug-and-play tracking adapter to enhance open-world foundation models (e.g., Grounding-DINO) for tracking arbitrary detected objects. NetTrack [40] introduced a plug-and-play framework that integrates a grounding model (e.g., Grounding-DINO[142]) to establish fine-grained object-text correspondences and a point tracker module (e.g., CoTracker[143]) to trace the trajectories of points of interest, thereby enhancing cross-frame object association performance.
Earlier RMOT approaches [144, 145, 146] mainly buit their systems based on a pretrained single-modal detection model (Deformable DETR), with cross-modal alignment and tracking modules integrated into this framework. Recently, iKUN [88] utilizes CLIP to encode the input query and each individual frame in video tubes, and subsequently employs a tracker to generate object trajectories. Furthermore, it introduces a knowledge unification module to guide the integration of visual features and textual information, producing text-specific visual representations and facilitating the selection of the most relevant tracking tubes. LaMOTer [42] conducts object tracking by detecting target objects in each frame with Grounding-DINO, predicting their locations in the next frame using an improved Kalman filter, and associating them to form the final tracks.
IV-B Video Segmentation
Open Vocabulary Video Instance Segmentation (OV-VIS) extends traditional VIS by enabling the segmentation and tracking of arbitrary object categories not seen during training, leveraging vision-language models to generalize beyond a fixed label set. In current research, the mainstream approach for this task leverages CLIP’s open-vocabulary capabilities by designing additional network modules appended after the CLIP encoder to achieve image-to-video transfer. OV2Seg [43] constitutes the inaugural work to incorporate CLIP into a post-network fine-tuning strategy for OVVIS. It employs a DETR-like architecture with object-centric queries to predict masks and perform cross-frame matching for tracking. The query embeddings, which fuse information from current and previous frames, are directly compared with CLIP-encoded text features to achieve open-vocabulary classification. InstFormer [97] introduces instance-guided attention into CLIP’s visual encoder, leveraging segmentation results as external cues to aggregate text-aligned instance features without compromising CLIP’s pretrained vision-language matching knowledge (using LoRA [141]). Furthermore, it feeds the aggregated instance tokens into the tracking module to match instance tokens between previous frames and the current frame. BriVIS [51] employs a frozen video instance segmentor to generate frame-level instance queries and integrates multi-scale CLIP visual features for alternating short-term and long-term post-network temporal modeling. To model instance dynamics, BriVIS introduces a Brownian bridge to constrain instance features and adopts a contrastive objective, ensuring that the bridge center remains close to its corresponding class text embedding while distancing itself from irrelevant class texts. OVFormer [52] employs video-level instance queries, as opposed to image-level queries, and aligns video-instance queries with frozen CLIP features via a cross-attention mechanism to achieve unified segmentation and tracking across frames. OV-DVIS++ [53] decomposes the task into segment-tracking-refinement stages, progressively modeling the problem from local to global contexts. CLIP-VIS [44] directly leverages CLIP’s image features to generate masks by combining predicted object scores and mask IoU. To alleviate the effect of historical noise, it first applies the Hungarian algorithm to identify the $K$ most relevant frames from memory, and then performs instance tracking.
Referring Video Object Segmentation (RVOS) aims to segment and track a target object instance across video frames based on given natural language description, which typically specifies the object’s appearance, motion, and relational cues. Grounded-SAM-2 [98] and AL-Ref-SAM-2 [99] follow the framework of finetuning with extra model, which attempt to employ GroundingDINO [142] to detect objects in individual frames, then use the grounding results as prompts for independent segmentation foundation models (e.g., SAM-2 [147]) to produce high-quality object masks. ReferDINO [72] proposes an end-to-end learning framework based on GroundingDINO, which first introduce a object-consistent temporal enhancer to capture the cross-frame object interaction. Then, it build a query pruning block on top of the pre-trained multimodality encoders in GroundingDINO, explicitly reducing the number of tokens for efficient inference. By combining the above modules and a mask decoder, ReferDINO succesfully empower the image-based foundation model with the ability of cross-modal temporal reasoning and segmentation.
IV-C Temporal Video Grounding

### v2

@@ II Overview of Image-Language
LLaVA [114] integrates visual understanding with natural language processing through an end-to-end trainable architecture. As Figure 3 (d) shows, LLaVA consists of a pre-trained vision encoder CLIP-ViT and an LLM for multi-modal processing. It employs a simple projection layer to align visual representations with the LLM’s embedding space, enabling cross-modality fusion. LLaVA undergoes two-stage training: first, feature alignment pretraining on $\mathord{\sim}$595K image-text pairs with a frozen vision encoder; second, instruction fine-tuning using $\mathord{\sim}$158K GPT-4-generated multi-modal instruction-response pairs to enhance reasoning and conversational capabilities. LLaVA-1.5 [112] extends LLaVA by replacing its image encoder with CLIP-ViT-L and vision-language projection layer with MLPs, which achieve better multi-modality understanding. Moreover, the model also introduces a multi-patch partitioning strategy, which improves the efficiency of processing high-resolution images and facilitates dynamic-resolution inputs. LLaVA-NeXT [113] quadruples the input image resolution to 672×672 and supports flexible aspect ratios. Using optimized visual instruction tuning on <<1M samples, it improves reasoning, OCR, and world knowledge, achieving SOTA results and surpassing Gemini Pro in some tasks while retaining scalability.
As illustrated in Figure 3(e), InternVL [25] is a large-scale vision-language foundation model with a vision encoder—InternViT-6B and a language middleware called QLLaMA, initialized from a multilingual LLaMA and augmented with cross-attention layers and learnable queries, with a total of 8B parameters. These components are aligned via a three-stage training strategy: starting with contrastive pre-training on 4.98B filtered image-text pairs from sources; followed by generative training on 1.03 billion high-quality captions using a combination of Image-Text Contrastive, Image-Text Matching, and Image-grounded Text Generation losses; and concluding with supervised fine-tuning on 4 million instruction samples for multimodal dialogue. Several extensions have been developed. InternVL2 [24] extends the framework to handle video and multimodal data. InternVL3 [216] demonstrates improved GUI understanding and spatial reasoning capabilities through a single-stage native multimodal pre-training.
Qwen-VL [3] is a vision-language foundation model trained via a multi-stage strategy (from pre-training to supervised fine-tuning) rather than simple visual encoder-LLM concatenation. As shown in Figure 3(f), its architecture includes a ViT-based visual encoder, a position-aware adapter, and the Qwen language model. The model is pre-trained on a massive dataset of over 1.4 billion image-text pairs and fine-tuned through a multi-stage training process, including large-scale weakly supervised and multi-task supervised fine-tuning. A key strength of Qwen-VL is its comprehensive multimodal capability, supporting multiple image inputs, multilingual conversations, and fine-grained recognition through its multi-scale training strategy and high-resolution processing (up to 448×448 pixels). Qwen-VL-Chat [3] enhances Qwen-VL with sophisticated multi-image dialogue capabilities, supporting complex interleaved vision-language tasks through improved alignment training and safety mechanisms. Qwen2-VL [172] further enhances visual reasoning, context handling, and multilingual performance, while Qwen2.5-VL [4] achieves significant breakthroughs through its upgraded vision encoder with dynamic high-resolution support (up to 1536×1536). It demonstrates superior performance on complex tasks like chart and document analysis, fine-grained VQA, long-context multimodal dialogue, and complex instruction following, achieving new SOTA across multiple vision-language benchmarks.
@@
Discussion: All the above ILFMs not only open up new opportunities for developing language-based image understanding models, but also set the stage for their further application in more complex video analysis tasks, where the integration of temporal information becomes essential. Among these ILFMs, MDETR and G-DINO are featured by their stronger abilities in capturing fine-grained visual–textual correspondences, since they are explicitly trained with grounding supervision to align object regions with textual phrases. In contrast, CLIP is more adept at handling coarse-grained visual–textual correspondences, due to its global-level contrastive pretraining objective. Besides, LLM-based models such as LLaVA, InternVL, and Qwen-VL empower ILFMs with strong instruction-following abilities, which is crucial for addressing more complex tasks such as reasoning and planning. Thus, they are expected to become mainstream for some advanced applications in future.
III Overview of Image-to-Video
Transfer Learning
In this section, we aim to meticulously construct an extensible taxonomy, in which any existing image-to-video transfer learning method can be classified into at least one technical paradigm. It first categorizes transfer methods into frozen-features and adapted-features paradigms depending on whether the features extracted by ILFMs remain unchanged or undergo adaptation. More specific sub-categories under these two root categories will be elaborated in the following. It is noted that we focus on transfer learning [138, 218, 136] involving parameter updates, while zero-shot learning [174, 19, 143] are out of scope.
@@ III-A Frozen Features
Figure 5: Methods and architectures of transferring pre-trained ILFM to video domain via frozen features.
By treating video as a sequence of image frames in temporal order, a considerable number of existing methods directly leverage pretrained ILFMs to extract frozen frame-level features, regarding them as guidance signal, initial or auxiliary inputs to conduct image-to-video transfer learning. Specifically, representative lines of research under such principle include knowledge distillation, post-network tuning, and side-tuning.
@@ III-A1 Knowledge Distillation
As shown in Figure 5(a), within the knowledge distillation paradigm, the common practice is to employ a pretrained ILFM as the teacher model to enhance the cross-modal alignment capabilities of the visual and textual encoders in the downstream video-text understanding models. Various distillation objectives can be defined to facilitate targeted image-to-video knowledge transfer between the teacher and student encoders, often working in conjunction with another downstream loss. For instance, the distillation paradigm has been widely adopted in early Open-Vocabulary Multi-Object Tracking (OV-MOT) methods [97, 103, 144, 95] to inherit open-vocabulary capabilities from ILFMs like CLIP. For better distillation transfer performance, one critical design point is on how to effectively balance the learning of temporal modeling knowledge from downstream data and strong cross-modal alignment knowledge from the pretrained ILFMs.
@@ III-A2 Post Network Tuning
Many methods achieve temporal modeling by additionally introducing a post network to process the frozen features obtained for each sampled frame (shown in Figure 5(b)). One straightforward scheme here is to construct the post network with temporal attention mechanism, enabling the information exchange across varied frames. However, this paradigm is usually limited in performance, as it is hard to directly capture dynamic temporal patterns from static features. To more effectively harness the visual potential of frozen large-scale image-text models, researchers employ a Q-Former [89] architecture, utilizing a set of learnable queries to interact with features from all video frames. Through this approach, queries can dynamically capture temporal dependencies and motion patterns in videos, rather than supplying image features with extra temporal cues. This not only reduces computational costs but also unlocks the potential of ILFMs. Beyond general temporal modeling, many methods have incorporated more task-specific temporal modeling strategies, such as Hungarian algorithm or Brownian bridge-based association. In summary, this paradigm relies on high-quality image-text features and carefully designed, task-specific temporal modeling architectures to achieve strong video understanding performance.
@@ III-A3 Side-Tuning
as illustrated in Figure 5(c), Side-tuning [204] often introduces a learnable side network in parallel to the main information paths of image/text encoders in the ILFMs. It receives the layer-wise latent features extracted by ILFM encoders as input, facilitating hierarchical spatial modeling and temporal refinement in a coarse-to-fine manner. In practice, the side network often adopts a lightweight design, incorporating functions like text-modulated spatial pooling and temporal modeling operations, etc. During training, gradients are effectively confined within the side-blocks, ensuring both parameter and memory efficiency. Recent studies [118, 207] show that the system can be further enhanced by connecting side-blocks with the outputs of CLIP, which achieves better results on several downstream video-text understanding tasks.
Discussion: Frozen-feature-based image-to-video transfer learning frameworks offer notable advantages in adapting image-based foundation models to video–language understanding tasks. They maintain high parameter efficiency by freezing pretrained architecture and weights, thereby mitigating catastrophic forgetting of previously acquired knowledge. However, their performance may be limited by the inherent representational capacity bottleneck of static image-level features and their imperfect alignment with video task requirements, particularly when generalizing to novel scenarios.
III-B Adapted Features
Instead of directly using frozen features, many approaches perform image-to-video transfer learning by explicitly adjusting the backbone model architecture and parameters, which we classify as adapted features. In the following, we will elaborate on the techniques, including Full Fine-Tuning, Partial Tuning, Fine-Tuning with Extra Models, Fine-Tuning with Adapter, Fine-Tuning with LoRA, and Prompt Tuning, whose architectures are presented in Figure 6, respectively.
Figure 6: Methods and architectures of transferring pre-trained image-text model to video domain via adapted features.
@@ III-B1 Full Fine-Tuning
It is straightforward to empower a pre-trained ILFM with the ability of video-text understanding by building a temporal block on top of each block of ILFM and then simply fine-tune all the parameters involved in the temporal modeling components together with the ILFMs [155, 195, 71, 108]. Although simple and effective, these approaches indeed are less applicable in practice as they need a large-scale, well-defined, and annotated data and a large amount of computing power in the model training stage.
@@ III-B2 Partial Tuning
Partial tuning strikes a balance between full fine-tuning and more parameter-efficient methods like adapters or prompt tuning. This strategy involves updating only a specific subset of the pre-trained ILFM’s parameters, while freezing the rest of the model parameters. Recently, several methods [14, 50, 205] have introduced temporal attention modules into the original spatial attention modules, enabling ILFMs to acquire temporal–spatial modeling capability. Such tuning strategies strike a favorable balance between performance and computational cost, while still relying heavily on the inherent spatial modeling capacity of the pretrained ILFMs, as they do not involve full model training.
@@ III-B4 Fine-Tuning with Adapter
Adapter-based methods offer a parameter-efficient solution to adapting ILFMs to video-language understanding tasks. This is achieved by inserting lightweight, trainable modules, called adapters, into intermediate layers of the frozen pre-trained models  [18, 190, 125]. While adapters are a general-purpose mechanism, their specific design is task-dependent. For video understanding, a common and effective design is to construct the adapter as a stack of temporal attention and cross-frame interaction layers, which are crucial for modeling the temporal dynamics in video data. By keeping the parameters frozen, it retains ILFMs’ visual-semantic understanding capabilities while efficiently learns video-specific temporal patterns through adapters. Recently, this technique has been particularly attractive in scenarios with limited video annotations or computational budgets, and it emerges as a prominent direction for addressing downstream tasks like VideoQA, TVG, and video captioning.
@@ III-B5 Fine-Tuning with LoRA
Low-Rank Adaptation (LoRA) [62] provides an alternative paradigm for ILFM adaptation. Instead of directly learning to update model weights $\Delta W$, LoRA approximates it with two small, trainable low-rank matrices, i.e., $\Delta W=BA$, where $B\in\mathbb{R}^{d\times r}$, $A\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$. This implicitly constrains the model’s exploration within a low-dimensional subspace, effectively mitigating the risk of overfitting, which is particularly beneficial when labeled video data is scarce. Thus, LoRA is widely used to efficiently adapt large-scale foundation models to various downstream video tasks. However, as LoRA cannot model temporal information itself, it requires external temporal modules or networks, such as those used in full fine-tuning or post-network tuning.
@@ III-B6 Prompt Tuning
Prompt tuning has recently emerged as a highly parameter-efficient alternative for adapting ILFMs to video-related tasks [37, 133]. By freezing the backbone and introducing a small set of learnable prompts, the model is guided to capture temporal cues without modifying its core parameters. A common design is to share a single set of prompts across all frames to enforce temporal consistency, or to propagate prompts sequentially across frames so that information accumulates and evolves with motion. These temporal variants enable the frozen ILFMs to move beyond static frame modeling and capture motion cues and cross-frame dependencies. This strategy achieves efficient spatio-temporal adaptation with low computational cost, though its reliance on frozen backbones may restrict fine-grained temporal reasoning compared to full or partial fine-tuning.
Discussion: The aforementioned image-to-video transfer learning methods primarily adapt existing image–text foundation models to address video–text understanding tasks through two main directions: enhancing the encoders of established video understanding models or incorporating additional learnable temporal modules into ILFMs. Notably, these approaches are not mutually exclusive. Existing studies have shown that their integration can yield complementary benefits and further improve transfer performance. For example, fine-tuning with auxiliary models can be seamlessly combined with parameter-efficient strategies such as LoRA, facilitating the development of more advanced and robust video understanding models.
@@ IV Fine-grained Video-Text Understanding
In the following, we review methods that adapt existing pre-trained ILFMs to address specific video-text understanding tasks. Based on the required granularity of understanding, we categorize the video-text understanding tasks into fine-grained and coarse-grained classes, providing a clearer structural understanding and enabling researchers to identify methods most relevant to their research domains. In the following, we first introduce fine-grained video-text understanding, which requires precise localization of spatial regions (e.g., bounding boxes or segmentations) or temporal intervals.
IV-A Open Vocabulary Multi-Object Tracking
Open Vocabulary Multi-Object Tracking (OV-MOT) aims to track all the objects described by input text in an open-vocabulary manner, rather than relying on a predefined, closed set of categories. It is a task extended from traditional MOT, in which the objects of interest are explicitly referred to with natural language descriptions.
@@ IV-A Open Vocabulary Multi-Object Tracking
Most of existing OV-MOT methods [98, 88, 86, 103, 97, 212, 99, 144] mainly follow a frozen feature paradigm to transfer useful knowledge from pre-trained CLIP and G-DINO models. For instance, OVTrack [97] and OVTrack+ [103] develop knowledge distillation blocks to enhance the RoI features extracted by ResNet50 and FPN with pre-trained CLIP, Based on OVTrack, VOVTrack [144] further introduce object appearance self-consistency and spatial-appearance mutual-consistency constraints to capture spatial relationships across frames, where object appearance is directly extracted with CLIP. SLAck [99] introduces a spatial-temporal object attention mechanism to jointly model semantic, location, and appearance cues, in which knowledge in CLIP is distilled to enhance the semantic head. GLATrack [86] employs a pre-trained CLIP model to extract semantic information of target objects from the reference frame, which is subsequently used to mitigate the impact of erroneous features in key frames. OVTR [88] employs a post-network fine-tuning strategy, wherein a Transformer-based dual-branch decoder is introduced to process frozen features extracted by CLIP. These methods rely on coarse-grained image-text correspondences captured by CLIP, which restricts the model’s ability to capture the open-world objects with high dynamicity. NetTrack [212] employs G-DINO [117]) to establish fine-grained object-text correspondences and a point tracker (e.g., CoTracker [75]) to trace trajectories of points of interest, hence enhancing cross-frame object association. MASA [98] introduces a plug-and-play tracking adapter t o enhance open-world foundation models (e.g., G-DINO) for tracking arbitrary detected objects.
IV-B Temporal Video Grounding